Solve. Identify all double roots.
x^4 - 25x^2 = 0

Answer by unlockmath(1688)

Hello,
This x^4 - 25x^2 = 0 can be rewritten as:
x^2(x^2-25)=0
This can be factored as:
x^2(x-5)(x+5)=0
Solve for x and we get:
x=0
x=-5
x=5
Make sense?
RJ
www.math-unlock.com